High Tech Inc. is a virtual store that stocks a variety of calculators in its warehouse. Customer orders are placed; the orders are picked and packaged; and then orders are shipped to the customers. A fixed-order-quantity inventory control system (FQS) helps monitor and control these SKUs. The following information is for one of the calculators that High Tech stocks, sells, and ships.

Average demand 9.5 calculators per week
Lead time 5 weeks
Order cost $15/order
Unit cost $10.00
Carrying charge rate 0.30
Number of weeks 52 weeks per year
Standard deviation of
weekly demand
3.25 calculators
SKU service level 95 percent
Current on-hand inventory 37 calculators
Scheduled receipts 19 calculators
Backorders 4 calculators
  1. What are the economic order quantity? Round your answer to the nearest whole number.

      calculators

  2. What are the total annual order and inventory-holding costs for the EOQ? Use the rounded EOQ value in your calculations. Do not round any other intermediate calculations. Round your answer to the nearest dollar.

    $   

  3. What is the reorder point without safety stock? Round your answer to the nearest whole number.

      calculators

  4. What is the reorder point with safety stock? Use Appendix A to determine z-value. Round z-value to two decimal places. Do not round any other intermediate calculations. Round your answer to the nearest whole number.

      calculators

  5. Based on the previous information, should a fixed-order quantity be placed? Consider the reorder point with safety stock. Round your answer to the nearest whole number.

    A fixed-order quantity -Select-shouldshould notItem 5  be placed because the inventory position of   calculators is -Select-higher thanlower thanequal toItem 7  the reorder point with safety stock.
Answer #1

EOQ = [ 2 x demand xorder cost / holdign cost per annum]1/2

= [ 2 x 52x9.5 x15/0.3 x10]1/2 = 70.28 =70

Ordering cost = 52x9.5 /70 x 15 =105.85

Holding cost = 70x3/2 =105

Reorder point = demand during lead time = 5 x9.5 =47.5 =48

Reorder point with safety stock = Demand during lead time + z x sd x ( lead time)1/2

= 48 + 3.25 x1.64x ( 5)1/2

= 60

Current inventory status = inventory on hand - backorders + receipts

= 37-4+19 = 52

Since the ROP is 60, the order should be placed, as the current balance is below 60.